OpenStudy (anonymous):

For this problem that I'm working, I got 9/24 which would be reduced to 3/8. The problem is x+1/8+-1/4. Is my answer correct?

OpenStudy (anonymous):

is that an equation? please clarify

OpenStudy (amistre64):

what is the +- stand for?

OpenStudy (amistre64):

might be a typo for =

OpenStudy (anonymous):

in which case x = -3/8

OpenStudy (amistre64):

8(x+1/8 = 1/4) 8x + 1 = 8/4 8x +1 = 2 8x +1-1 = 2-1 8x = 1 8x/8 = 1/8 x = 1/8 if I read it correctly...

OpenStudy (amistre64):

if its a -1/4; then yeah, prifk is right :)

OpenStudy (anonymous):

until then.. we're in the dark..

OpenStudy (amistre64):

lol.... light some candles nd eat some PBJs :)